OpenStudy (anonymous):

Advanced Function Problem. For what values of C is the Fahrenheit temperature greater than the equivalent Celsius temperature?

OpenStudy (anonymous):

\[f=\frac{ 9 }{ 5 }c+32\] \[c=(f-32)\frac{ 5 }{ 9 }\]

OpenStudy (aum):

"For what values of C is the Fahrenheit temperature greater than the equivalent Celsius temperature?" \[ f > c \\ \frac{ 9 }{ 5 }c+32 > c \\ \text{Solve for c.} \]

OpenStudy (anonymous):

why did you make f=c? i dont get that part.

OpenStudy (aum):

I did not put f = c anywhere.

OpenStudy (anonymous):

Instead f you put c is less than the equation for Celsius.

OpenStudy (anonymous):

oh wait nvm

OpenStudy (anonymous):

sorry about that got little confused. with the equations

OpenStudy (anonymous):

not confused mixed up with the equation i should say.

OpenStudy (aum):

Can you solve the above inequality for c?

OpenStudy (anonymous):

yup im trying that.

OpenStudy (anonymous):

\[\frac{ 5 }{ 9 }c+32>c\] \[32>c-\frac{ 5 }{ 9 }c\] \[32>\frac{ 4 }{ 9 }c\] \[\frac{ 32(9) }{ 4 }>\frac{ 4c }{ 4 }\] \[72>c\]

OpenStudy (anonymous):

in the textbook it says c is greater than -40 for the answer what am i doing wrong?

OpenStudy (anonymous):

got it nvm found my mistake.

OpenStudy (aum):

9/5 c +32 > c 9/5 c - c > -32 4/5 c > -32 4c > -160 c > -40